OpenStudy (princess_shadow_sob):

WILL MEDAL IF YOU HELP A proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (5, 4) is on the graphed line. Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points)

OpenStudy (princess_shadow_sob):

@pooja195

OpenStudy (princess_shadow_sob):

@benlindquist

OpenStudy (princess_shadow_sob):

@benlindquist

OpenStudy (princess_shadow_sob):

@benlindquist

OpenStudy (princess_shadow_sob):

@benlindquist

OpenStudy (princess_shadow_sob):

@benlindquist

OpenStudy (benlindquist):

ook

OpenStudy (princess_shadow_sob):

oh and it didnt come with graph

OpenStudy (benlindquist):

ok let me get one

OpenStudy (princess_shadow_sob):

i just said it DIDNT come with one

OpenStudy (benlindquist):

I'm getting one myself

OpenStudy (princess_shadow_sob):

ok

OpenStudy (princess_shadow_sob):

hurry

OpenStudy (benlindquist):

test?

OpenStudy (benlindquist):

proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (5, 4) is on the graphed line. Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points)

OpenStudy (benlindquist):

X=price in dollars. Y=graphed. and the ordered pairs are 5,4

OpenStudy (princess_shadow_sob):

ya i said that in my question can u help

OpenStudy (benlindquist):

ill try, i think i know

OpenStudy (benlindquist):

@mathmale

OpenStudy (benlindquist):

@PokemonGirlNerd11

OpenStudy (princess_shadow_sob):

@dakid88

OpenStudy (princess_shadow_sob):

@dakid88

OpenStudy (dakid88):

I assume the relationship is linear in the form y = mx +c If (5,4) is on the line we can calculate the equation gradient m = change in y over change in x, = 4/5 y = 4/5 * x +c 4 = 4/5 * 5 +c c = 0 y = 4/5 x Part a if x = 1 then y = 4/5 part b is 2 times (5,4) so 10 pounds of potatoes cost 8 dollers

OpenStudy (princess_shadow_sob):

thx ^_^

OpenStudy (dakid88):

welcome dont forget to medal @Princess_Shadow_SOB